# Gemstones: A Model Suite for Multi-Faceted Scaling Laws

### v1

@@ 3.1 Optimal Learning Rates for Gemstones
Figure 4: Approach 1 prescriptions. Row one: Validation loss over FLOPs (left) and GPU hours (right). We use Approach 1 to find the optimal points on the convex hull in each setting, marked with black crosses. Row two: We fit a line to the tokens per parameter of empirically optimal models and find a slightly higher, but still constant, tokens per parameter prescription than Hoffmann et al. (2022).  Hoffmann et al. (2022)’s Approach 1 creates $250$ logarithmically-spaced FLOPs bins per order of magnitude, and in red we plot the minimizers over these bins, and the scaling law fitted to these minimizers (binning). Clearly, their Approach 1 is not well-suited for our data, and our convex hull approach is better when we select fewer models to fit our law on. Extended plot in Figure 9.
3.2 Fitting Scaling Laws
We fit scaling laws using methods similar to approach $1$ and $3$ from Chinchilla (Hoffmann et al., 2022). We fit all laws using the log perplexity of all trained models on a sample of $100$ million tokens from a fixed, held-out validation set from the training distribution. We also collect log perplexity values for a range of open source models (Team et al., 2024a, b; Touvron et al., 2023; Dubey et al., 2024; Yang et al., 2024a, b) on the same validation data to allow for a comparison between our predictions and a selection of widely used models. We design a specialized FLOP counting function as we find that simple rules of thumb (e.g., FLOPs= $6\times parameters$ (Hoffmann et al., 2022)) do not accurately account for differences in FLOPs between extremely wide and narrow architectures. We discuss this further and present our function in Appendix G.
Following prior work, we plot the Epoch AI Replication (Besiroglu et al., 2024) of Chinchilla (Hoffmann et al., 2022) on all plots and use the coefficients for Kaplan plotted by Porian et al. (2024) which were extracted from the original paper (Kaplan et al., 2020).
A More Robust Approach to Fitting Compute-Optimal Laws. The first approach in Hoffmann et al. (2022) fits a scaling law by plotting the loss against FLOPs for a range of architectures, and then fitting a line to the pareto-optimal architecture for each FLOP count (see Figure 4). Following Hoffmann et al. (2022), we refer to this as “Approach 1”. As we use a constant learning rate, we can use all recorded validation losses to fit our law. Hoffmann et al. (2022) and Kaplan et al. (2020) select model shapes so densely that they have a near-optimal architecture at each FLOP count. This works when all architectures lie in a 1D space (parameterized by parameter count), as each model is optimal in some FLOP regime, and the lower envelope is densely populated. In our two dimensional exploration (varying width and depth), some models are never optimal, and the ones that are do not densely populate the envelope. We therefore need to propose a new fitting method for less data-dense regimes.
Our New Method: The Convex Hull. We fit a lower convex hull to our loss curves. This hull is only supported by a sparse set of optimal models. This naturally excludes sub-optimal models that lie above the convex hull of optimality, and as we will show, this makes the resulting scaling law far more robust to model selection choices.
Why We Skip Approach 2. Another method to fit scaling laws is to put model runs into isoFLOP bins and choose the best parameter count in each bin.  Hoffmann et al. (2022) call this “Approach 2”. Our 2-dimensional set of models do not finely cluster into isoFLOP bins, meaning our data is not easily amenable to Approach $2$, hence we exclude this approach from our analysis.  Hu et al. (2024) also eschew this approach.
Figure 5: Approach 3 laws with the parametrization shown in Equation 2. We see the prescribed optimal width-depth ratio increases with the FLOPs (left) budget and the optimal tokens per parameter decreases as the FLOPs budget increases (right). We see slight bumpiness in the lines due to the integer constraints we enforce on the attention heads, we also plot with this constraint removed in Figure 16.
Prescribing Optimal Widths and Depths by Fitting Power Laws. The final approach described by Hoffmann et al. (2022) is to fit a parametric formula to the loss values with the ansatz
$L(p,T)=\frac{A}{p^{\alpha}}+\frac{B}{T^{\beta}}+\varepsilon$ (1)
where $p$ is parameter count and $T$ is tokens. We fit our models using L-BGFS (Liu & Nocedal, 1989) with a Huber loss ($\delta=10^{-4}$) between the empirical log loss and the model prediction, and use multiple initializations following Besiroglu et al. (2024). We ablate to check that our fitting procedure is robust to the size of the grid of initializations and the choice of delta in Section D.5.
Our broad selection of model architectures enables us to study scaling laws that predict loss as a function of not only parameter count, but also model aspect ratio. For this purpose, we consider a perturbation of the standard scaling law with additional terms to account for the impact of model width and depth.
$L(w,d,p,T)=\frac{A}{w^{\alpha}}+\frac{B}{d^{\beta}}+\frac{C}{p^{\gamma}}+\frac{D}{T^{\zeta}}+\varepsilon.$ (2)
Here, $w$ is the hidden dimension of the model, $d$ is the number of layers. All of $A,\alpha,B,\beta,C,\gamma,D,\zeta$ and $\varepsilon$ are optimized to fit our scaling runs. We choose this form as it predicts the same behaviors as the standard law if width and depth are not implicated in any systematic trends. We ablate another possible form of this law in Section D.6.
4 Experiments
In Section 4.1 we use our new convex hull fitting method to make a scaling law for the compute-optimal tokens-to-parameters ratio, and our new power law approach to provide a prescription for the compute-optimal width-to-depth ratio. We show how many seemingly innocuous design choices such as the learning rate schedule can significantly impact these prescribed scaling laws in Section 4.2. Finally, we reflect on the benefits of training compute-optimal and time-optimal models in Section 4.3.
@@ 4 Experiments
Figure 6: We demonstrate the variability in fitting scaling laws by resampling our data many different ways. We label prescriptions found using Approach 1 with “Approach $1$” in the legend, otherwise approach $3$ is used. All tokens counts available are used to fit the laws unless stated otherwise in the legend, for example $\leq 100B$ means that only token counts less than or equal to $100B$ are used in fitting.
@@
Chinchilla Reduced Sampling: We subsample our data to be as close as possible to the token counts and model sizes that Hoffmann et al. (2022) use to fit their scaling laws and also fit new scaling laws on this subset of Hoffmann et al. (2022) data. Details in Section 4.2.
4.1 Sizing Up Our Scaling Laws Against Prior Laws and Industry Models
Approach 1.  In Figure 4 (row one), we see our validation losses plotted as both a function of FLOPs (left) and GPU hours (right) for the first $100$ billion tokens of training. We calculate GPU hours from the average recorded optimizer step time for each model.
In Figure 4 (row two), we see the prescription of the fitted laws for tokens per parameter. We see that the tokens per parameter prescription of our Approach 1 fitting is also close to constant, like Hoffmann et al. (2022), but slightly higher, suggesting more tokens should be used per parameter in the model. We extend this plot showing predicted total parameters, tokens, and over multiple ablations in Appendix C. We give a more detailed plot of each model’s individual validation loss in Appendix F.
Approach 3.  This uses a parametric function to predict the loss when given parameter count and number of tokens. As our data is intentionally designed to cover a wide variety of widths and depths, we can also predict the optimal aspect ratio of a model for a given FLOP budget. We do this by optimizing over the law shown in Equation 2 with four terms so we can explicitly optimize the width and depth.
We find that convex optimization struggles to find optimal points due to the integer constraint we impose on the number of attention heads in the model. Instead, we use a brute force approach by selecting approximately $10^{8}$ model shapes and tokens counts, then use the law to predict the loss for each model, and choose the minimal loss among those in each FLOP range. In Figure 5 we plot the output of this process as a solid line labeled “Approach 3 (Ours)”. In the left of the figure, we see that the prescribed width-depth ratio increases with the FLOP budget, but that width-depth ratio increases relatively slowly even as FLOPs is increased by many orders of magnitude (corroborating certain observations in Levine et al. (2020)). We also plot a selection of widely used models trained by industry estimating their cumulative FLOP expenditures (x-axis values) using FLOPs/token = $6\times parameters$ to calculate FLOPs with their published training token counts. In Figure 5 (right), we see that our optimal tokens per parameter ratio tracks the prescription of Kaplan et al. (2020) more closely than Hoffmann et al. (2022); the prescribed tokens per parameter decreases as the number of available FLOPs increases.
4.2 A Rainbow of Scaling Laws
To demonstrate the sensitivity of scaling laws to design choices, we fit laws with various assumptions and model selection rules. We begin by fitting laws of the classical form (without width-depth), and use equation $4$ from Hoffmann et al. (2022) to provide compute-optimal parameter count prescriptions. In Figure 6 we show the optimal predictions of multiple possible laws fitted on different data subsets, visualizing the amount of variation possible under an array of configurations.
One particular dimension of variability we wish to highlight briefly here is the interplay between model selection and the derived law. To do this, we select 5 models from Gemstones that have an analogous model in Hoffmann et al. (2022) (using data extracted by Besiroglu et al. (2024)) with similar parameter count and aspect ratio. We select Gemstones checkpoints with token counts nearly matching the Hoffman points. We call this “Chinchilla Reduced Sampling.” We then fit scaling laws to both of these sub-sampled datasets. We find that fitting Hoffmann’s data using this reduced sampling results in an increased slope relative to fitting on all data. Meanwhile this subsampling reduces the slope of the line fit on Gemstones. This highlights that scaling law fitting can be quite sensitive to seemingly innocuous changes in model selection for both the Gemstones and the simpler model family selected by Hoffman. Notably, there are $5$ models in this subset for both Hoffmann et al. (2022) and our data, this meets the rule of thumb given by Choshen et al. (2024) for the minimum number of models should be used to fit a scaling law.
We also perform many other ablations in Figure 6, and we summarize these here. First, we fit laws both including and excluding embedding parameter count, which both Pearce & Song (2024) and Porian et al. (2024) find to be a primary explanation of the discrepancies between the prescriptions found by Kaplan et al. (2020) and Hoffmann et al. (2022). We also show the impact of fitting on our cooldown and learning rate ablation datasets in turn, seeing that both choices have a noticeable impact on the prescription for optimal parameter count. Next, we remove checkpoints from our data to simulate having only trained for $100$ billion tokens or only having data for token counts greater than $120$ billion. Finally, we also fit to only models with a width of $512$ to isolate the role of depth. Fitting on only these models creates a drastic difference in the law, significantly reducing the predicted optimal rate of growth of parameters as a function of FLOPs, highlighting the need for including diverse model shapes in the fitting data.
4.3 The Price of Stepping Off the Scaling Law
@@
Figure 7: The inefficiency of training models with suboptimal widths and depths. We plot the FLOPs (left) and GPU Hours (right) overspend after training our Gemstones for $300$ billion tokens. We define the overspend as how many resources (FLOPs or GPU hours) are required for a model with a given width-depth configuration to reach some target loss, relative to the models that achieve that target loss the fastest (the “points on (pareto)-frontier”). We can see that the skinny models (top-left, dark points) use many more FLOPs or GPU hours to reach a target loss than the wide models. We note that these inefficiencies exist in our training setup because we only use tensor parallelism and not pipeline parallelism.
In the previous sections we provide prescriptions for the optimal width, depth, parameters, and tokens as a function of total training FLOPs. As we have seen, industry models don’t always comply with our prescriptions. At the same time, the variability in our analyses suggests that scaling laws are inherently fuzzy, providing only rough estimates of optimal settings. With this in mind, we ask a natural question: if one happens to choose a sub-optimal point in design space, how bad can it really be?
@@
By analyzing the cost of stepping off of the scaling law, we find that some kinds of design errors are more damaging than others. In particular, one pays a steep price for training models that are too narrow, and it is better to err on the side of too wide. We also find that training on more tokens than is strictly recommended (aka “overtraining”) is typically quite efficient in terms of pushing down loss.
@@
If You Value Your Time, Train Wide Models. We first show that in our training setup, training wider models is far more efficient than training deep models. In Figure 7, we reflect on the consequences of suboptimal architectural choices, by considering how much of a given resource—FLOPs or GPU hours—would be “overspent” to reach any target loss value with the plotted architecture rather than the prescribed width and depth. We find that choosing to train “skinny” models (top left) wastes many FLOPs and GPU hours. The scale of overspend is quite different however, with the least efficient models only overspending about $50\%$ on FLOPs but wasting more than $200\%$ of the GPU hours spent by the best configuration. In other words, in the time taken to train a single (very) suboptimal model to the desired loss value, one could train three optimal-width-depth models. We note that while the time-optimal models tend to be the wider ones, this is probably due to our training scheme. Similar to other open-source efforts such as OLMo et al. (2024), we do not make any use of pipeline parallelism, and only employ tensor parallelism (using a hybrid data and tensor parallel algorithm similar to the ubiquitous Fully Sharded Data Parallel strategy). In summary, for standard parallelism implementations, wider models are simply easier to scale, but as a result our observations regarding resource overspending may not generalize to other parallelism strategies.
Figure 8: Quantifying the cost of overtraining. We simulate deviations from our prescriptions to assess their impact on model performance by increasing the optimal token count prescribed by Equation 2 by an overtraining factor. We then optimize the model shape to achieve the lowest loss possible at each FLOP budget and overtraining factor. Note that $10^{0}$, or 1×1$\times$1 ×, is the prescribed optimal point. We take four FLOP budgets (title of each plot) and plot the loss as a function of overtraining factor and see that under or overtraining increases predicted loss but by only a small amount. We plot the losses of selected open source models on our validation set to help ground the y-axis ranges.
Scaling Laws Predict That Overtraining Is Efficient.
@@
In Figure 5, we see the optimal predictions from our law of the form shown in Equation 2. We can shift those optimal points to simulate overtraining. To do this, we fix a FLOP budget and trace out a path of model sizes and corresponding token counts to remain within that budget. For each model size and token count, we record the “overtraining factor,” which is the selected number of training tokens divided by the optimal number of tokens for that model shape. An overtraining factor of less than one corresponds to undertraining the model, and a factor greater than one represents overtraining. We show the results of this process in Figure 8. We see that overtraining does increase predicted loss at a given FLOP count but that these curves are actually quite flat. We include the loss values of open source models on our own validation set to allow readers to contextualize the y-axis values. Especially at high FLOP counts, overtraining becomes quite efficient in that it results in fairly small elevations in loss for a relatively large reduction in model size.

### v2

@@ A.1 Optimal Learning Rates for Gemstones
A.2 Software and Data
We train all models using a fork of litgpt [AI, 2023] enhanced with AxoNN [Singh and Bhatele, 2022, Singh et al., 2024] tensor parallelism. We open source all models used in our analysis to Hugging Face [Wolf et al., 2020] and the logging from training on Weights and Biases in json format.
Appendix B Extended Related Works
Scaling laws are broadly applied to many areas of machine learning, such as machine translation. Ghorbani et al. [2021] split the parameters term into two, one for each of encoder and decoder, and similarly to Gordon et al. [2021] analyze the relationship between BLEU scores and scaling laws. Zhang et al. [2022] and Bansal et al. [2022] study the impact of architecture choice on the scaling law, finding increasing data or parameters can compensate for worse architectural decisions.
There are multiple works analyzing whether scaling laws can be used to predict downstream performance. Ruan et al. [2024] show that scaling laws can be predictive of benchmark performance. Caballero et al. [2023] observe that traditional scaling laws cannot capture complex behaviors like non-monotonic trends nor inflection points. They propose broken scaling laws–a piecewise or smoothly broken power law–that better predicts performance of both upstream and downstream tasks.
Scaling laws have also been applied to sparse architectures. Clark et al. [2022] analyze how the number of experts can be used in the law, studying both linear and quadratic interactions for many types of routing models. Yun et al. [2024] extend this, analyzing the trade offs between optimal training and optimal inference. Krajewski et al. [2024] find that with optimal settings a Mixture of Experts model always outperforms a transformer model at any computational budget. Frantar et al. [2023] focus on weight sparsity within foundation models, adding a multiplicative parameter on the parameters term in the law. Model shape has also been analyzed for sparse mixture of expert models and in the context of finetuning. Krajewski et al. [2024] use the ratio between the standard feed-forward hidden dimension and the hidden dimension of an individual expert to allow their law for mixture of expert models to predict the width of the experts.
These techniques are not limited to generative text modeling only; they have also been applied to multi-model models. Henighan et al. [2020] find optimal model size can be described as a power law for model modeling including images and video. The authors also find that model size does not help ‘strong generalization’ for problem solving. Aghajanyan et al. [2023] analyze text, images, code and speech, presenting a scaling law to describe the competition between these modalities and describe a regime for optimal hyperparameter transfer from the unimodal to multimodal regimes. Liang et al. [2024] look at scaling laws for diffusion transformer models. Li et al. [2024a] analyze scaling laws for vision encoder commonly used to encode image inputs for transformer model backbones, finding increasing the size of the encoder alone can lead to performance degradation in some cases.
Further analyses using scaling laws have extended to analyzing finetuning and data limited scaling. Hernandez et al. [2021] find that finetuning is much more compute efficient when the pretraining ignored. Zhang et al. [2024] study parameter efficient finetuning regimes find a multiplicative law is better for the finetuning setting than the classical additive law used by others. Muennighoff et al. [2023] analyze the data constrained training regimes, finding epoching data up to four times is as good as training on deduplicated data in terms of reducing loss.
Zero-shot Hyperparameter Transfer
The ability to train a series of models with extremely different parameter counts is an implicit requirement of any scaling law analysis. Work on zero-shot hyperparameter transfer across transformer model widths is mature [Yang et al., 2021, Everett et al., 2024, Hayou and Yang, 2023, Dey et al., 2024]. Achieving transfer across diverse model depths is less well studied, especially in transformer language models [Bordelon et al., 2024, Yang and Hu, 2021, Yang et al., 2023]. While Yang et al. [2023] argue that depth transfer requires scaling in $1/\sqrt{L}$ because this is the unique regime with maximum feature diversity, more recently Dey et al. [2025] argue that instead the scaling should be in $1/L$.
B.1 Chinchilla Equation 4
Hoffmann et al. [2022] use the variable names $N$ and $D$ for the number of parameters and number of tokens respectively, defining their parameterized form as:
$\hat{L}(N,D)\triangleq E+\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}},$ (6)
Equation 4 of Hoffmann et al. [2022] is defined as:
$N_{opt}(C)=G{\left(\frac{C}{6}\right)}^{a},\quad D_{opt}(C)=G^{-1}{\left(\frac{C}{6}\right)}^{b},$ (7)
 $\quad\text{ where }\quad G={\left(\frac{\alpha A}{\beta B}\right)}^{\frac{1}{\alpha+\beta}},\quad a=\frac{\beta}{\alpha+\beta},\text{ and }b=\frac{\alpha}{\alpha+\beta}.$ (8)
Figure 10: All possible model shapes we could have chosen based on our architecture within $\pm 5\%$ are shown as circles. The points we selected are highlighted as stars, including the two extra points we select to have four models of width $512$.
Appendix C Data Sampling
We plot the entire space of all possible models subject to our design constraints discussed in Figure˜10. While exploring the impact of finer grained depth differences during our experiments, we decided to add two additional models slightly outside the $\pm 5\%$ tolerance band at the 100100M scale; for $width=512$, in addition to the originally chosen depths of $12$ and $13$, we added $11$ and $14$; these appear as a dense collection of $4$ points at the same widthwidth.
Appendix D Extrapolation to Larger Models
To quantify the robustness of our scaling laws to changes in model scale, we perform two types of extrapolation analyses. In the first experiment, we hold out the “2B2B” parameter models, fit scaling laws to the samller models, and then test the extrapolation performance of the fitted law. Due to the $10\%$ tolerance margin in our parameter count strata, in actuality we fit benchmark scaling laws using models with less than $1.8B$ parameters and extrapolate to models with more than $1.8B$ parameters. In Figure˜12 we see that extrapolating to predict error offers a much better fit than we see in Figure˜12 when predicting accuracy.
Figure 11: Benchmark Scaling Law for Error. We fit a law of the form shown in Equation˜3 to benchmark results sampled at every $10$ billion tokens using models with less than $1.8B$ parameters and observe a tight fit when extrapolating to models with more than $1.8B$ parameters.
Figure 12: Benchmark Scaling Law for Accuracy. We fit a law of the form shown in Equation˜5 to benchmark results sampled at every $10$ billion tokens for ARC, HellaSwag and MMLU using models with less than $1.8B$ parameters and observe a poor fit when extrapolating to models with more than $1.8B$ parameters.
In a second type of experiment, we again analyze extrapolation as a function of model size but this time quantifying prescription robustness in terms of estimated versus actual validation loss (using approach 3). Following Choshen et al. [2024], we report the mean absolute relative error (ARE) over the last $30\%$ of training tokens ($250b$ to $350b$ tokens for Gemstone models). We find the ARE when fitting on all checkpoints and then testing on models with more than $1.8B$ parameters is $0.68\%$. When only fitting on models with less than $1.8B$ parameters and extrapolating to models with more than $1.8B$ parameters, the ARE is $0.63\%$. Choshen et al. [2024] find ARE’s of up to $4\%$ are typically used to distinguish between modeling choices, implying that our extrapolation error is well within the acceptable range.
Appendix E Leave-One-model-Out Analysis
To evaluate how robust our scaling laws are to a different aspect of our experimental design, we estimate the variability in our fitting process caused by our model selection using a leave-one-out analysis. In Figure˜13 we re-fit the same scaling law multiple times leaving each one of the model shapes out in turn using both Approach 1 and Approach 3. We visualize these results by plotting the minimum and maximum tokens per parameter ratios yielded across all leave-one-out trials along with the prescription based on all data (bounds of the gray shaded region vs the green line). While the implications of the precise min/max values are somewhat up to interpretation, compared to the difference between our fit, Kaplan’s, and Chinchilla’s, the relative narrowness of the gray region suggests that any disagreement in tokens per parameter prescription we find versus prior work is unlikely to simply be an artifact of (our) specific model selection. Finally, comparing the left and right sides of Figure˜13, the smaller gray region in the former suggests that Approach 1 is less sensitive to this model re-sampling process than Approach 3. We hypothesize that at least some of this increased robustness in fit can be attributed to our novel variance-reducing convex hull method applied in Approach 1.
Figure 13: Leave-One-Out Validation. We leave each model out and refit the law to obtain the gray shaded region (min/max ratio across all trials) and compare it with the law fit to all data in green. Left) shows the result when Approach 1 is used, and right) shows the result of using Approach 3.
Appendix F Variability in fitting
In Table˜2, we show a similar table to Table˜1 but exclude embeddings from the parameter count in the fitting process. We also visualize all fitted lines in Figure˜14.
Table 2: We demonstrate the variability in fitting scaling laws by resampling our data many different ways. The slope can be viewed as the exponent in the power law relationship $parameters=constant\cdot compute^{exponent}$. Grouping by fitting approach and choice to include embeddings, in the final column ‘Delta’ we show the change in slope produced by the ablations against the corresponding base law fit on the full set of hot data. Values with an absolute magnitude greater than 0.05 are highlighted in orange, and those exceeding 0.1 are highlighted in red. We see that the reduced sampling has a large impact on the slope of the law and that Approach 1 is more sensitive than Approach 3. We plot these prescriptions in Figure˜14 and show this table with embeddings included in the parameter count in Table˜1.
Tokens Cooldown LR Ablation Embeddings Slope Delta
Kaplan et al. [2020] 0.7300
Approach 1 (no Embeds)
all 0.5689
$\leq 100b$ 0.6269 0.0579
>120b>120b 0.9666 0.3977
all 0.6224 0.0535
all 0.7242 0.1552
Approach 3 (no Embeds)
all 0.7141
$\leq 100b$ 0.7030 -0.0111
>120b>120b 0.7350 0.0209
all 0.6929 -0.0211
all 0.7104 -0.0037
@@ Appendix F Variability in fitting
Figure 14: We demonstrate the variability in fitting scaling laws by resampling our data many different ways. We label prescriptions found using Approach 1 with “Approach 11” in the legend, otherwise approach 33 is used. All tokens counts available are used to fit the laws unless stated otherwise in the legend, for example $\leq 100B$ means that only token counts less than or equal to $100B$ are used in fitting.
@@
Chinchilla Reduced Sampling: We subsample our data to be as close as possible to the token counts and model sizes that Hoffmann et al. [2022] use to fit their scaling laws and also fit new scaling laws on this subset of Hoffmann et al. [2022] data. Details in Section˜4.2.
Appendix G Individual Benchmark Results
In Figure˜15, we see zero-shot MMLU scores of our larger models are quite non-trivial at $28\%-33\%$, despite being trained on an open dataset, without a cooldown period, or any sort of post-training.
Figure 15: Individual benchmark performance. We benchmark all models using the 200, 250, 300, and 350B token checkpoints, converting this to FLOPs using the formula shown in Appendix˜L. Models show increasing accuracy with depth when constrained to approximately the same FLOP budget (vertically aligned points) on many benchmarks. We plot the average benchmark accuracy in Figure˜8.
Appendix H The Price of Stepping Off the Scaling Law
@@
Figure 16: The inefficiency of training models with suboptimal widths and depths. We plot the FLOPs (left) and GPU Hours (right) overspend after training our Gemstones for $300$ billion tokens. We define the overspend as how many resources (FLOPs or GPU hours) are required for a model with a given width-depth configuration to reach some target loss, relative to the models that achieve that target loss the fastest (the “points on (pareto)-frontier”). We can see that the skinny models (top-left, dark points) use many more FLOPs or GPU hours to reach a target loss than the wide models. We note that these inefficiencies exist in our training setup because we only use tensor parallelism and not pipeline parallelism but highlight how to easily transfer these results to other environments in Section˜4.4.
@@
By analyzing the cost of stepping off of the scaling law, we find that some kinds of design errors are more damaging than others. We also find that training on more tokens than is strictly recommended (aka “overtraining”) is typically quite efficient in terms of pushing down loss.
@@
If You Value Your Time, Train Wide Models. We first show that in our training setup, training wider models is far more efficient than training deep models. In Figure˜16, we reflect on the consequences of suboptimal architectural choices, by considering how much of a given resource—FLOPs or GPU hours—would be “overspent” to reach any target loss value with the plotted architecture rather than the prescribed width and depth. We find that choosing to train “skinny” models (top left) wastes many FLOPs and GPU hours. The scale of overspend is quite different however, with the least efficient models only overspending about $50\%$ on FLOPs but wasting more than $200\%$ of the GPU hours spent by the best configuration. In other words, in the time taken to train a single (very) suboptimal model to the desired loss value, one could train three optimal-width-depth models. We note that while the time-optimal models tend to be the wider ones, this is probably due to our training scheme. Similar to other open-source efforts such as OLMo et al. [2024], we do not make any use of pipeline parallelism. In summary, for standard parallelism implementations, wider models are simply easier to scale, but as a result our observations regarding resource overspending may not generalize to other parallelism strategies.
Appendix I Scaling Laws Predict That Overtraining Is Efficient.
@@
Similarly to Gadre et al. [2024], we can shift optimal points to simulate overtraining. To do this, we fix a FLOP budget and trace out a path of model sizes and corresponding token counts to remain within that budget. For each model size and token count, we record the “overtraining factor,” which is the selected number of training tokens divided by the optimal number of tokens for that model shape. An overtraining factor of less than one corresponds to undertraining the model, and a factor greater than one represents overtraining. We show the results of this process in Figure˜17. We see that overtraining does increase predicted loss at a given FLOP count but that these curves are actually quite flat. We include the loss values of open source models on our own validation set to allow readers to contextualize the y-axis values. Especially at high FLOP counts, our laws predict overtraining becomes quite efficient in that it results in fairly small elevations in loss for a relatively large reduction in model size.